OpenStudy (anonymous):

Simplify the radical expression Square root of 45? 1.) 5 square root of 3 2.) 3 square root of 5 3.) 3 square root of 15 4.) 15

OpenStudy (anonymous):

3 square root of 5

OpenStudy (anonymous):

Thanks how did you get that?

OpenStudy (er.mohd.amir):

45=3*3*5 sqrt(45)=sqrt(3*3*5)=3*sqrt(5)

OpenStudy (anonymous):

Thats correct